OpenStudy (anonymous):

Compute the derivative y=6sqrtx

OpenStudy (anonymous):

\[y=6\sqrt{x}\]

OpenStudy (cwrw238):

rewrite it as 6 x^0.5 now use the formula if y = ax^n y' = an x^(n-1)

OpenStudy (cwrw238):

lets try an easy one: y = 3x^3 y' = 3*3 x^(3-1) = 9 x^2

OpenStudy (cwrw238):

does that help?

OpenStudy (anonymous):

yes so it would be y=6x^6 Right ? then 6*6x^(6-1) 36x^5 is that right?

OpenStudy (cwrw238):

no 6 x^0.5 compare this with a x^n a = 6 and n = 0.5 ok?

OpenStudy (cwrw238):

so derivative is an x^(n-1) a=6,n=0.5 so derivative = 6 * 0.5 x^(0.5 - 1) = 3x ^(-0.5)

OpenStudy (anonymous):

ok but how did you get n

OpenStudy (cwrw238):

i think i confused you with 3x^3: using 2 3's instead of something like 3 x^2

OpenStudy (cwrw238):